Put the following function: y=x^2-6x+8 in the form y=a(x-h)2=k
Any advice on how to find this would be great! I tried to follow the instructions given by the book but it did not work out so well! It says to find the answer by completing the square. I tend to get lost in that process when you add the 8 to both sides.

Answer by stanbon(75887)

Put the following function: y=x^2-6x+8 in the form y=a(x-h)2+k
----------------
Rewrite as:
y-8+? = x^2-6x+?
To complete the square you need (-6/2)^2=9
Put that in place of the ? to get:
y-8+9=x^2-6x+9
Rewrite as:
y+1 = (x-3)^2
Then:
y=(x-3)^2-1
Cheers,
Stan H.